# Collaborative-Online-Learning-Enabled Distributionally Robust Motion Control for Multi-Robot Systems

### v1

@@ Appendix C Proof of Lemma 11
We note that the two conditions in $\widetilde{\mathbb{D}}^{n}$ (47) also have the corresponding equivalent forms as in (101) and (102). Therefore, for the first condition in $\widetilde{\mathbb{D}}^{n}$, we verify its form corresponding to (101d) below.
@@
where the generalized inequality (a) follows from the definitions of $\overline{\gamma}^{n}$ and $\breve{\gamma}^{n}$ in (52), while the generalized inequality (b) arises from $p^{j}\in\mathbb{D}^{j}$. Therefore, $\widetilde{p}^{n}$ satisfy the first condition in $\widetilde{\mathbb{D}}^{n}$. For the second condition in $\widetilde{\mathbb{D}}_{\ell,t,k}^{n}$, we verify its form corresponding to (102b), which requires verifying the following inequality.
@@
$\overset{(b)}{\preceq}$ $\sum_{j\in\mathbb{M}_{n}}\frac{\gamma^{j}}{\widetilde{\gamma}^{n}}\left\{(\varepsilon^{j}+2\beta^{j})\hat{\Sigma}^{j}+3(\hat{\mu}^{j}-\hat{\widetilde{\mu}}^{n})(\hat{\mu}^{j}-\hat{\widetilde{\mu}}^{n})^{T}\right\}+3\widetilde{\beta}^{n}\hat{\widetilde{\Sigma}}^{n}$
$\overset{(c)}{\preceq}$ $\sum_{j\in\mathbb{M}_{n}}\bar{\gamma}_{\varphi}^{n}\left\{(\varepsilon^{j}+2\beta^{j})\hat{\Sigma}^{j}+3(\hat{\mu}^{j}-\hat{\widetilde{\mu}}^{n})(\hat{\mu}^{j}-\hat{\widetilde{\mu}}^{n})^{T}\right\}+3\widetilde{\beta}^{n}\hat{\widetilde{\Sigma}}^{n}$
@@
where the generalized inequality (b) arises from $p^{j}\in\mathbb{D}_{\ell,t,k}^{j}$, and the generalized inequality (a) follows from the definitions of $\bar{\gamma}^{n}$. Therefore, $\widetilde{p}^{n}$ satisfies the second condition in $\widetilde{\mathbb{D}}^{n}$. This completes the proof that $\widetilde{p}^{n}\in\widetilde{\mathbb{D}}^{n}$.
@@ Appendix D Proof of Theorem 14
$s.t.$ $\left(\begin{array}[]{cc}\Lambda^{n}&\xi^{n}\\ (\xi^{n})^{T}&\tau^{n}\end{array}\right)\succeq 0\qquad\Omega^{n}\succeq 0$ (151)
@@ Appendix E Proof of Theorem 16
We claim that, due to the spatial allocation protocol, if (11) is feasible at time $t$, the collision between the two robots will never occur. Next, we analyze the collision avoidance between robots and obstacles. According to Theorem 8, Lemma 9 and 11, we know that
@@
At each control time $t$, for the obstacle $\ell\in\mathbb{O}_{i,t}$, the probability of collision avoidance at the predicted time $t+1$ is given by the DR-CVaR constraint (10) as $\alpha^{u}$. As a result, the control trajectory $y_{i}(t+1)=y_{i}(1|t)$ is collision avoidance with obstacle $\ell$ with at least $\alpha_{\ell}\alpha^{u}$. For the obstacle $\ell\notin\mathbb{O}_{i,t}$, according to Section 5.2, the formulation of the avoidance constraints with obstacle $\ell$ is either based on the ambiguity set from a past time step or based on the support set. Therefore, the trajectory also avoids the collision with obstacle $\ell$ with at last a probability of $\alpha_{\ell}\alpha^{u}$. In summary, the total collision avoidance probability of $y_{i}(t+1)=y_{i}(1|t)$ is at least $\prod_{\ell=1}^{L}\alpha_{\ell}(\alpha)^{L}$.
Based on the above analysis, the single-step trajectory $y_{i}(t+1)=y_{i}(1|t)$ has a collision avoidance probability of at least $\prod_{\ell=1}^{L}\alpha_{\ell}(\alpha^{u})^{L}$. Therefore, if (11) is feasible at each time t=0,,T1t=0,...,T-1, the trajectory $y_{i}(1),...,y_{i}(T)$ is collision-free with at least $\prod_{\ell=1}^{L}(\alpha_{\ell})^{T}(\alpha^{u})^{TL}$. This completes the proof. $\quad\square$
@@ Appendix F Proof of Theorem 20
(1) Constraints (11b)-(11f). Due to the construction of (160) and the assumption of the terminal set $\mathcal{X}_{i}^{f}$, the feasibility of candidate solution (160) is straightforward.
(2) Constraint (11g). (i) For the case k=1,,K2k=1,...,K-2, we can verify the feasibility of the candidate solution (160) with respect to the newly computed separating hyperplane $\mathcal{H}_{i,j,t+1,k}^{rob}$. If it is infeasible, the separating hyperplane $\mathcal{H}_{i,j,t,k+1}^{rob}$ at time $t$ can be directly used to maintain feasibility. (ii) For the case k=K1k=K-1, we know that $x_{i}(K-1|t+1)=x_{i}(K|t)\in\mathcal{X}_{i}^{f}(t+K)$ and $x_{j}(K-1|t+1)=x_{j}(K|t)\in\mathcal{X}_{j}^{f}(t+K)$. According to the assumption of the terminal set $\mathcal{X}_{i}^{f}$, we can obtain $(y_{i}(K|t)\oplus\mathcal{R}_{i})\cap(y_{j}(K|t)\oplus\mathcal{R}_{j})=\emptyset$. Due to the convexity of $\mathcal{R}_{i}$ and $\mathcal{R}_{j}$, a separating hyperplane $\mathcal{H}_{i,j,t+1,K-1}^{rob}$ can always be found to maintain feasibility.
(3) Constraint (11h). We first consider about the obstacle $\ell\in\mathbb{O}_{i,t}\land\ell\in\mathbb{O}_{i,t+1}$. (i) For the case k=1,,K2k=1,...,K-2, due to the feasibility at time $t$, we can obtain
$\hskip-2.0pty_{i}(k|t+1)^{T}h_{i,\ell,t,k+1}^{obs}+\mathcal{S}_{\mathcal{R}_{i}}(h_{i,\ell,t,k+1}^{obs})+g_{i,\ell,t,k+1}^{obs}+r$ (161)
$=y_{i}^{*}(k+1|t)^{T}h_{i,\ell,t,k+1}^{obs}+\mathcal{S}_{\mathcal{R}_{i}}(h_{i,\ell,t,k+1}^{obs})+g_{i,\ell,t,k+1}^{obs}+r\leq$ $0$
where $g_{i,\ell,t,k+1}^{obs}$ is calculated as follows.
$g_{i,\ell,t,k+1}^{obs}=$ (162)
$\mathop{\sup}\limits_{\upsilon\in\widetilde{\mathbb{D}}_{\ell,t,k+1}}{\rm CVaR}_{\alpha^{u}}^{\upsilon}(-\varphi^{T}h_{i,\ell,t,k+1}^{obs}+\mathcal{S}_{\mathcal{O}_{\ell}}(-h_{i,\ell,t,k+1}^{obs}))\leq 0$
The same normal vector as at time $t$ is used at time $t+1$, i.e., $h_{i,\ell,t+1,k}^{obs}=h_{i,\ell,t,k+1}^{obs}$, and the bias $g_{i,\ell,t,k+1}^{obs}$ is computed as follows.
$g_{i,\ell,t+1,k}^{obs}=$ (163)
$\mathop{\sup}\limits_{\upsilon\in\mathbb{D}_{\ell,t+1,k}}{\rm CVaR}_{\alpha^{u}}^{\upsilon}(-\varphi^{T}h_{i,\ell,t,k+1}^{obs}+\mathcal{S}_{\mathcal{O}_{\ell}}(-h_{i,\ell,t,k+1}^{obs}))\leq 0$
According to the assumption $\mathbb{D}_{\ell,t+1,k}\subseteq\mathbb{D}_{\ell,t,k+1}$ and $\mathbb{D}_{\ell,t,k+1}\subseteq\widetilde{\mathbb{D}}_{\ell,t,k+1}$ through Lemma 11, we can obtain $\mathbb{D}_{\ell,t+1,k}\subseteq\widetilde{\mathbb{D}}_{\ell,t,k+1}$ and $g_{i,\ell,t+1,k}^{obs}\leq g_{i,\ell,t,k+1}^{obs}$. Therefore, if (161) is satisfied, constraint (11h) based on $h_{i,\ell,t+1,k}^{obs}$ and $g_{i,\ell,t+1,k}^{obs}$ obtained above is also satisfied. In other words, we can always employ the separating hyperplane $\mathcal{H}_{i,\ell,t,k+1}^{obs}$ at time $t$ as a safety alternative to both maintain feasibility and satisfy the DR-CVaR constraint (10). (ii) For the case k=K1k=K-1, we know that $x_{i}(K-1|t+1)=x_{i}(K|t)\in\mathcal{X}_{i}^{f}(t+K)$. According to the assumption of the terminal set $\mathcal{X}_{i}^{f}$, we can obtain $(y_{i}(K|t)\oplus\mathcal{R}_{i})\cap(\mathbb{W}_{\ell,t+K}\oplus\mathcal{O}_{\ell})=\emptyset$. Due to the convexity of $\mathcal{R}_{i}$, $\mathcal{O}_{\ell}$, and $\mathbb{W}_{\ell,t+K}$, a separating hyperplane $\mathcal{H}_{i,\ell,t+1,K-1}^{obs}$ can always be found to maintain feasibility.
We then consider about the obstacle $\ell\notin\mathbb{O}_{i,t+1}$. According to the construction of $\mathcal{H}_{i,\ell,t,k}$ in Section 5.2, we can assume, without loss of generality, that $\mathcal{H}_{i,\ell,t,k}$ for k=1,,Kl1k=1,...,K-l-1 is obtained based on the compressed ambiguity set $\widetilde{\mathbb{D}}_{\ell,t-l,k+l}$, and $\mathcal{H}_{i,\ell,t,k}$ for k=Kl,,K1k=K-l,...,K-1 is obtained based on the support set $\mathbb{W}_{\ell,t+k}$. In particular, when $l=0$, it represents the case $\ell\in\mathbb{O}_{i,t}$. We then analyze the feasibility at time $t+1$. (i) For k=1,,Kl2k=1,...,K-l-2, the separating hyperplane $\mathcal{H}_{i,\ell,t+1,k}^{obs}$ is calculated based on $\widetilde{\mathbb{D}}_{\ell,t-l,k+l+1}$, and thus setting $\mathcal{H}_{i,\ell,t+1,k}^{obs}=\mathcal{H}_{i,\ell,t,k+1}^{obs}$ can maintain feasibility. (ii) For k=Kl1,,K2k=K-l-1,...,K-2, the separating hyperplane $\mathcal{H}_{i,\ell,t+1,k}^{obs}$ is calculated based on $\mathbb{W}_{\ell,t+k+1}$, and thus setting $\mathcal{H}_{i,\ell,t+1,k}^{obs}=\mathcal{H}_{i,\ell,t,k+1}^{obs}$ can also maintain feasibility. (iii) For k=K1k=K-1, we can always obtain $\mathcal{H}_{i,\ell,t+1,K-1}^{obs}$ based on the assumption of $\mathcal{X}_{i}^{f}$ as discussed before.
We finally consider about the obstacle $\ell\notin\mathbb{O}_{i,t}\land\ell\in\mathbb{O}_{i,t+1}$. As we discussed before, we can assume that $\mathcal{H}_{i,\ell,t,k}$ for k=1,,Kl1k=1,...,K-l-1 is obtained based on $\widetilde{\mathbb{D}}_{\ell,t-l,k+l}$, and $\mathcal{H}_{i,\ell,t,k}$ for k=Kl,,K1k=K-l,...,K-1 is obtained based on $\mathbb{W}_{\ell,t+k}$. We discuss the feasibility at time $t+1$ below. (i) For k=1,,Kl2k=1,...,K-l-2, due to the assumption $\mathbb{D}_{\ell,t+1,k}\subseteq\mathbb{D}_{\ell,t,k+1}$, we can obtain $\mathbb{D}_{\ell,t+1,k}\subseteq\mathbb{D}_{\ell,t-l,k+l+1}\subseteq\widetilde{\mathbb{D}}_{\ell,t-l,k+l+1}$. Therefore, we can reduce this discussion to the case where $\ell\in\mathbb{O}_{i,t}\land\ell\in\mathbb{O}_{i,t+1}$, and conclude that $\mathcal{H}_{i,\ell,t,k+1}^{obs}$ can serve as a safety alternative to both maintain feasibility and satisfy the DR-CVaR constraint (10) at time $t+1$. (ii) For k=Kl1,,K2k=K-l-1,...,K-2, we note that the separating hyperplane $\mathcal{H}_{i,\ell,t,k+1}^{obs}$ is calculated based on the support set $\mathbb{W}_{\ell,t+k+1}$, thus $\mathcal{H}_{i,\ell,t,k+1}^{obs}$ inherently satisfies the DR-CVaR constraint (10) at time $t+1$. Therefore, setting $\mathcal{H}_{i,\ell,t+1,k}^{obs}=\mathcal{H}_{i,\ell,t,k+1}^{obs}$ can maintain feasibility. (iii) For k=K1k=K-1, we can always obtain $\mathcal{H}_{i,\ell,t+1,k}^{obs}$ based on the assumption of $\mathcal{X}_{i}^{f}$ as discussed before.
(4) Constraint (11i). Based on the definition of $\mathcal{X}_{i}^{f}$, (11i) is evidently feasible for the candidate solution (160).

### v2

@@ Appendix C Proof of Lemma 11
We note that the two conditions in $\widetilde{\mathbb{D}}^{n}$ (48) also have the corresponding forms as in (102) and (103). Therefore, for the first condition in $\widetilde{\mathbb{D}}^{n}$, we verify its form corresponding to (102d) below.
@@
where the generalized inequality (a) follows from the definitions of $\overline{\gamma}^{n}$ and $\breve{\gamma}^{n}$ in (53), while the generalized inequality (b) arises from $p^{j}\in\mathbb{D}^{j}$. Therefore, $\widetilde{p}^{n}$ satisfy the first condition in $\widetilde{\mathbb{D}}^{n}$. For the second condition in $\widetilde{\mathbb{D}}_{\ell,t,k}^{n}$, we verify the corresponding sufficient condition in (103a), which requires verifying the following inequality.
@@
$\overset{(b)}{\preceq}$ $\sum_{j\in\mathbb{M}_{n}}\frac{\gamma^{j}}{\widetilde{\gamma}^{n}}\left\{(\varepsilon^{j}+3\beta^{j})\hat{\Sigma}^{j}+3(\hat{\mu}^{j}-\hat{\widetilde{\mu}}^{n})(\hat{\mu}^{j}-\hat{\widetilde{\mu}}^{n})^{T}\right\}+3\widetilde{\beta}^{n}\hat{\widetilde{\Sigma}}^{n}$
$\overset{(c)}{\preceq}$ $\sum_{j\in\mathbb{M}_{n}}\bar{\gamma}_{\varphi}^{n}\left\{(\varepsilon^{j}+3\beta^{j})\hat{\Sigma}^{j}+3(\hat{\mu}^{j}-\hat{\widetilde{\mu}}^{n})(\hat{\mu}^{j}-\hat{\widetilde{\mu}}^{n})^{T}\right\}+3\widetilde{\beta}^{n}\hat{\widetilde{\Sigma}}^{n}$
@@
where the generalized inequality (b) arises from $p^{j}\in\mathbb{D}_{\ell,t,k}^{j}$, and the generalized inequality (c) follows from the definitions of $\bar{\gamma}^{n}$. Therefore, $\widetilde{p}^{n}$ satisfies the second condition in $\widetilde{\mathbb{D}}^{n}$. This completes the proof that $\widetilde{p}^{n}\in\widetilde{\mathbb{D}}^{n}$.
@@ Appendix D Proof of Theorem 14
$s.t.$ $\left(\begin{array}[]{cc}\Lambda^{n}&\xi^{n}\\ (\xi^{n})^{T}&\tau^{n}\end{array}\right)\succeq 0\hskip 18.49988pt\Omega^{n}\succeq 0$ (152)
@@ Appendix E Proof of Theorem 17
Due to the spatial allocation protocol, if (12) is feasible at $t$, the collision between two robots will never occur. Next, we analyze the collision avoidance between robots and obstacles. According to Theorem 8, Lemma 9 and 11, we know that
@@
At each control time $t$, for the obstacle $\ell\in\mathbb{O}_{i,t}$, the probability of collision avoidance at time $t+1$ is given by (11) as $\alpha^{u}$. As a result, the control trajectory $y_{i}(t+1)=y_{i}(1|t)$ is collision avoidance with obstacle $\ell$ with at least $\alpha_{\ell}\alpha^{u}$. For the obstacle $\ell\notin\mathbb{O}_{i,t}$, according to Section 5.3, the formulation of the avoidance constraints with obstacle $\ell$ is either based on the ambiguity set from a past time step or based on the support set. Therefore, the trajectory also avoids the collision with obstacle $\ell$ with at last a probability of $\alpha_{\ell}\alpha^{u}$. In summary, the total collision avoidance probability of $y_{i}(t+1)=y_{i}(1|t)$ is at least $\prod_{\ell=1}^{L}(\alpha_{\ell}\alpha)^{L}$. Therefore, if (12) is feasible at each time t=0,,T1t=0,...,T-1, the trajectory $y_{i}(1),...,y_{i}(T)$ is collision-free with at least $\prod_{\ell=1}^{L}(\alpha_{\ell}\alpha^{u})^{T}$. This completes the proof. $\quad\square$
@@ Appendix F Proof of Theorem 20
First, for constraints (12b)-(12f), by the construction of the candidate solution (161) and the definition of the terminal set $\mathcal{X}_{i}^{f}$, the feasibility of candidate solution (161) is straightforward.
Second, for the collision avoidance affine constraints (12g) and (12h), the safe update rules for separating hyperplanes proposed in Section 5.3 ensure that, at each time step, the adopted separating hyperplanes are compatible with the current committed trajectory (i.e., the candidate output). As a result, these affine constraints remain feasible for the candidate solution (161).
Finally, for the terminal constraint (12i), by the construction of the terminal set $\mathcal{X}_{i}^{f}$ in Definition 16, the terminal state corresponding to the candidate solution satisfies $\hat{x}_{i}(K|t+1)\in\mathcal{X}_{i}^{f}$. Therefore, the terminal constraint is also satisfied.